# YSO Jets Driven by Magnetic Pressure Generated through Stellar Magnetosphere–Disk Interaction

### v1

@@ 3 Result
In this section, we present an overview of our simulation results, highlighting key features of the jet structure. Figure 1 shows a 3D rendering of the jet in our model, where blue and red contours represent the outflow in the upper and lower hemispheres, respectively. Each hemisphere contains two layers: the outer layer encloses gas moving at speeds exceeding ±$\pm 100$ km s-1, while the inner layer highlights gas with speeds exceeding ±$\pm 200$ km s-1, illustrating the presence of a powerful bipolar jet. To avoid obscuring the jet, the disk is visualized as a toroidal structure around the midplane, with most of its mass removed for clarity. The white lines represent magnetic fieldlines, and velocity streamlines are shown by the black lines.
@@
The partial opening of the magnetosphere during the early stages of simulations involving a stellar dipole is a well-known process (e.g. A. P. Goodson & R. M. Winglee, 1999; M. M. Romanova et al., 2009; P. S. Lii et al., 2014). In this phase, the initially closed outer dipole field lines that thread the disk inflate and eject poloidal magnetic loops. This magnetic ejection results in the partial opening of the magnetosphere as the system undergoes its initial adjustment.
@@
The initial adjustment results in a characteristic magnetic field geometry. To illustrate the magnetic field geometry and its relationship with the outflow, we show in Figure 2 the magnetic fieldlines overplotted on the projected $\hat{z}$-direction velocity $v_{z}^{p}$, defined as
@@
so outflow in both the upper and lower hemispheres is positive whereas inflow is negative. The magnetic field geometry can be crudely divided into four zones: the permanently closed stellar magnetosphere near the star (enclosed by the magenta line in Figure 2), the $-\hat{z}$ direction opened stellar field attached to the star (roughly enclosed by the cyan lines), the large-scale open $+\hat{z}$-direction field (similar to the original disk field, the outer region marked by the green lines), and the less ordered field lines lying in between the former two (between the green and cyan lines). This overall magnetic field structure remains consistent throughout the simulation after the initial adjustment. Since the magnetic field configuration plays a crucial role in launching the outflow, we first examine its properties before discussing each outflow component in detail.
@@
Figure 2: Magnetic field geometry, represented by the magnetic field lines on a meridian slice, after the initial adjustment in the simulation. The background color map shows the projected $\hat{z}$-direction velocity $v_{z}^{p}$ (equ. 9). The magnetic field structure can be roughly divided into four distinct zones, delineated by the magenta, cyan, and green lines. The magenta line encloses the permanently closed stellar magnetosphere; the cyan line marks the boundary of opened stellar field lines that still thread the star; the green line outlines the extent of the large-scale magnetic field lines threading the disk. The region between the cyan and green lines corresponds to a reconnecting, outflow-driving zone where the interaction between stellar and disk fields powers the outflow.
3.1 Magnetic flux evolution
To quantify the opening of the magnetosphere field during simulation, Figure 3 shows the downward magnetic flux ($-\hat{z}$ direction) at $z=\pm 5~{}\mathrm{au}$ as a function of time for the upper and lower hemispheres, respectively. Since the disk’s magnetic field is oriented upward, any downward flux at these heights must originate from the opening of the stellar magnetosphere. For reference, Figure 3 also shows the total stellar magnetospheric flux, defined as the total downward magnetic flux at the midplane. Due to magnetic reconnection of the innermost magnetospheric loops inside the star, the total stellar magnetic flux decreases slightly at the beginning of the simulation before stabilizing at approximately $-8\times 10^{26}~{}\mathrm{G\ cm^{2}}$. The opened magnetospheric flux remains around $-8\times 10^{25}~{}\mathrm{G\ cm^{2}}$, accounting for about 10% of the total magnetospheric flux.
@@ 3.1 Magnetic flux evolution
Although the newly opened field lines within $1~{}\mathrm{au}$ originate from the stellar magnetosphere rather than the original disk field, they thread the disk vertically and link the dense midplane to the less dense surface layers, just as the original disk field. This connection facilitates the launching of a slower disk wind at larger cylindrical radii, which we examine in the following subsection.
@@
3.2 Disk Wind from Opened Magnetospheric Field
The disk wind is a slower, high-density outflow at large cylindrical radius compared to the jet (see § 3.4 below). To illustrate the properties and launching mechanism of the disk wind, we show in Figure 4 the azimuthally averaged quantities in the disk wind region. Panels (a) and (b) show the density and the projected $\hat{z}$-direction velocity, respectively. Any outflow away from the disk midplane has a positive value. The disk wind has a density $\approx 10^{-13}~{}\mathrm{g\ cm^{-3}}$ near the base (around $r\approx 0.5~{}\mathrm{au}$), and a density of $\approx 10^{-14}~{}\mathrm{g\ cm^{-3}}$ at higher altitudes. The outflow velocity of the disk wind is around $10^{6}~{}\mathrm{cm\ s^{-1}}$, slower than the jet but still above the escape velocity at high altitudes.
Y. Tu et al. (2025a) identify two components for the disk wind: an MRI-active disk wind and a laminar disk wind. In their case, the inner disk wind is MRI-unstable, leading to a turbulent outflow, whereas the outer disk wind is MRI-stable, resulting in a laminar wind. The MRI stability is measured using the dimensionless parameter (Y. Tu et al., 2025a)
@@ 3.2 Disk Wind from Opened Magnetospheric Field
where $H=R/10$ is the disk scale height and $\lambda$ is the most unstable MRI wavelength given by (Z. Zhu & J. M. Stone, 2018)
$\lambda=2\pi\sqrt{\frac{16}{15}}\frac{v_{A}}{\Omega}.$ (11)
Here, $v_{A}$ is the Alfvén speed and $\Omega$ is the local orbital frequency. Following Y. Tu et al. (2025a), we choose the range $100.1<q<$ 10 to denote where MRI is most efficient.
Figure 4(c) shows this $q$ parameter, illustrating a zone between the jet and the disk wind in both hemispheres where $q$ is within the optimal range. The disk wind can thus be divided into the MRI-active disk wind at a small cylindrical radius and the MRI-stable disk wind at a larger cylindrical radius, particularly along the upper hemisphere. However, similar to Y. Tu et al. (2025a), the MRI-active disk wind eventually merges with the laminar disk wind at high altitudes, where MRI can no longer be sustained. As a result, at these altitudes, the primary distinction is between the jet and the disk wind.
@@
measuring the ratios between the thermal and ram pressures and the magnetic pressure. The plasma-$\beta<1$ in the disk wind indicates that thermal energy is negligible compared to magnetic energy. Kinetic-$\beta>1$ in the disk indicates that kinetic energy (primarily from rotation) dominates the magnetic energy.
@@
While the magnetic force generally points in the $-\hat{R}$ direction (towards the axis), its magnitude is insufficient to counteract the outward centrifugal force. This imbalance leads to an increasing opening angle of the disk wind.
@@
and following Y. Tu et al. (2025a), we define the projected effective magnetic pressure and tension forces in cylindrical coordinates by removing the $B_{z}\frac{dB_{z}}{dz}$ term that does not contribute to overall acceleration:
@@
It is evident that magnetic pressure drives the disk wind outflow, not the magnetic tension force. This result is consistent with the geometric argument and conclusion in Y. Tu et al. (2025a) that only magnetic pressure can drive an outflow when open magnetic field lines thread the disk.
Figure 4: Overview of the outflow, particularly the disk wind at a large cylindrical radius surrounding the jet. Panel (a) and (b) show the density and projected $\hat{z}$-direction velocity ($v_{z}^{p}$, equ. 9) respectively; panel (c) shows the normalized MRI-turbulence wavelength (equ. 10; panel (d) and (e) show the plasma-$\beta$ and (equ. 12) kinetic-$\beta$ (equ. 13) respectively. The lower panels compare different force components in the disk wind. Panel (f) shows the ratio between the cylindrical-$\hat{R}$ direction magnetic force and centrifugal force. Panel (g), (h), (i), and (j) show the magnitude of the gas acceleration due to magnetic forces. The quantities are total gas acceleration in cylindrical-$\hat{R}$ direction, total magnetic acceleration in $\hat{z}$- irection, the effective magnetic pressure component (equ. 17), and the effective magnetic tension component (equ. 18) respectively. These disk wind properties and forces show that the disk wind in this model is very similar to the disk wind in Y. Tu et al. (2025a).
3.3 Opened Stellar Field
As discussed in Sec. 3.1, part of the stellar magnetospheric field opens up. Half of the opened field threads the disk and becomes a “disk field”; the other half remains attached to the star and becomes an opened stellar field.
To illustrate the geometry of the opened stellar magnetic field, Figure 5(a) presents a density slice through the meridional plane, overlaid with magnetic field lines. On the displayed scale, the magnetic field in the polar regions above and below the star points in the $-\hat{z}$ direction, opposite to the field threading the disk (e.g., Figure 4), as expected.
@@ 3.3 Opened Stellar Field
Although the majority of the open stellar field does not directly contribute to jet launching, it serves as a structural backbone for the jet in the magnetically dominated region. As we demonstrate in Section 3.4, the jet is primarily driven by toroidal magnetic pressure. However, a magnetic tower consisting predominantly of toroidal loops is inherently susceptible to kink instabilities (R. Moll et al., 2008; M. Huarte-Espinosa et al., 2012). The strong open stellar magnetic field in the polar region, which accounts for most of the open stellar flux, helps stabilize the jet by threading these loops along the polar axis, mitigating instabilities.
@@
Figure 5: A zoom-in view focusing on the permanently closed stellar magnetosphere (around the midplane) and the opened stellar magnetic field threading the star (in the polar region above and below the star). Panel (a) shows the real density traced by the “passive scalars”, highlighting the magnetospheric accretion streams. The sphere in the middle is the inner boundary $r_{\mathrm{fix}}$, where a star resides (see sec. 2 for a more detailed description of the setup). Panels (b) and (c) show the $B_{z}$ field strength and $v_{z}^{p}$ (equ. 9) respectively, illustrating the magnetic and dynamical properties of the region close to the star. Panels (d) and (e) show the plasma-$\beta$ (equ. 12) and kinetic-$\beta$ (equ. 13) respectively, indicating the magnetosphere and the polar cavity are both magnetically dominated.
3.4 Magnetospheric jet
@@
The jet is launched from the interaction between the star, its magnetosphere, and the surrounding disk, and propagates in the low-density polar cavity. Figure 6 outlines the jet’s location and key characteristics. At lower altitudes ($\lesssim 0.5~{}\mathrm{au}$), the jet is confined between the opened stellar field near the polar axis (approximately delineated by the magenta line) and the disk wind/atmosphere at larger cylindrical radii (indicated by the green dashed line). At higher altitudes, the jet expands to fill the polar cavity, as shown in Figure 4[b].
To assess the relative contributions of various energy components in the jet-launching region—namely, gas thermal energy, gas kinetic energy, and magnetic energy—we examine the plasma-$\beta$ and kinetic-$\beta$ in Figure 6(c) and (d). As expected, the plasma-$\beta$ remains well below unity throughout the jet region, except in the disk midplane, indicating that magnetic energy dominates over thermal energy. The kinetic-$\beta$ is also less than 1 inside the jet-launching polar cavity, but rises above 1 just outside the jet boundary. This transition is a key feature in the launching process and is analyzed in more detail in Section 4.
@@
Since both plasma-$\beta$ and kinetic-$\beta$ are below order unity in the jet-launching polar cavity, we focus on the magnetic forces to establish a foundation for our detailed examination of the jet-launching mechanism in section  4. The lower panels of Figure 6 illustrate the magnetic forces in the jet region. Figure 6[e] shows the total magnetic force, revealing that the acceleration of the outflow occurs primarily along the surface of the vertically extended disk, which also marks the boundary of the jet-launching cavity. Above approximately 1 au in both hemispheres, acceleration becomes less coherent (with alternating regions of acceleration and deceleration), indicating that the bulk of the flow acceleration has largely completed beyond this altitude. This is confirmed by velocities in the $\hat{z}$ direction reaching several $10^{7}~{}\mathrm{cm~{}s^{-1}}$ below 1 au, with minimal subsequent increase in velocity above this height.
@@
As demonstrated in Y. Tu et al. (2025a), in systems with ordered, large-scale magnetic fields threading the disk, only magnetic pressure forces, not magnetic tension forces, can launch outflows in global models. Our simulations confirm this finding. Figure 6(f) and (g) display the effective magnetic tension acceleration (equ. 18) and effective magnetic pressure acceleration (equ. 17), respectively. Magnetic tension acceleration generally points toward the midplane, while magnetic pressure acceleration points away from it. More specifically, Figure 6(h) shows the acceleration due to toroidal magnetic pressure gradient, defined as
@@
The toroidal magnetic pressure accounts for most of the magnetic pressure-driven jet acceleration. In the following section, we will examine the mechanism generating this crucial toroidal magnetic field (Section 4).
Figure 6: Overview of the jet. Upper panels focus on the magneto-hydrodynamic properties of the jet; lower panels focus on the magnetic forces. Panels (a) and (b) show the density and projected $\hat{z}$-direction ($v_{z}^{p}$, equ. 9) respectively, showing the jet lies in the low-density polar cavity. Panels (c) and (d) show the plasma-$\beta$ (equ. 12) and kinetic-$\beta$ (equ. 13) respectively, showing the jet is a magnetically-dominated structure. Panels (e), (f), (g), and (h) show the total magnetic force, the effective magnetic tension force (equ. 18), the effective magnetic pressure force (equ. 17), and the toroidal magnetic pressure force (equ. 19) respectively.
4 Jet launching mechanism: “Load-fire-reload”
This section presents a detailed description and analysis of the jet-launching mechanism in our model. The jet is launched through a cyclic process comprising three distinct phases: load, fire, and reload. Each phase corresponds to characteristic changes in the magnetic field configuration and dynamics. In the following subsections, we describe each step in detail and analyze how they collectively drive the jet-launching behavior observed in the simulation.
4.1 Load: charging $B_{\phi}$ by star-disk differential rotation

### v2

@@ III Results
Figure 1 shows a 3D rendering of the jet in our model, where blue and red contours represent the outflow in the upper and lower hemispheres, respectively. Each hemisphere contains two layers: the outer layer encloses gas moving at speeds exceeding $\pm 100$ km s-1, while the inner layer highlights gas with speeds exceeding $\pm 200$ km s-1, illustrating the presence of a powerful bipolar jet. To avoid obscuring the jet, the disk is visualized as a toroidal structure around the midplane, with most of its mass removed for clarity. The white lines represent magnetic fieldlines, and velocity streamlines are shown by the black lines.
@@
The partial opening of the magnetosphere during the early stages of simulations involving a stellar dipole is a well-known process (e.g. Goodson and Winglee, 1999; Romanova et al., 2009; Zanni and Ferreira, 2013; Lii et al., 2014; Ireland et al., 2021). In this phase, the initially closed outer dipole field lines that thread the disk inflate and eject poloidal magnetic loops. This magnetic ejection results in the partial opening of the magnetosphere as the system undergoes its initial adjustment.
@@
The initial adjustment results in a characteristic magnetic field geometry. To illustrate the magnetic field geometry and its relationship with the outflow, we show in Figure 2 the time- and azimuthal-averaged magnetic fieldlines overplotted on the projected $\hat{z}-$direction velocity $v_{z}^{p}$, defined as
@@
so outflow in both the upper and lower hemispheres is positive whereas inflow is negative. Since the magnetic field connecting the star and the disk is essential in driving the outflow, we provide a sketch of the key elements involved in jet-launching in Figure 2(b). The magnetic field geometry can be crudely divided into four zones: the permanently closed stellar magnetosphere near the star (enclosed by the magenta line in Figure 2), the $\hat{z}$ direction opened stellar field attached to the star (roughly enclosed by the cyan lines), the large-scale open $\hat{z}-$direction field (similar to the original disk field, the outer region marked by the green lines), and the less ordered field lines lying in between the former two (between the green and cyan lines). This overall magnetic field structure remains consistent throughout the simulation after the initial adjustment. Since the magnetic field configuration plays a crucial role in launching the outflow, we first examine its properties before discussing each outflow component in detail.
@@
Figure 2: Left panel: Magnetic field geometry, represented by the time- and azimuthal-averaged magnetic field lines during the later stages of the simulation. The background color map shows the projected $\hat{z}-$direction velocity $v_{z}^{p}$ (equ. 9). Right panel: a cartoon illustration of the magnetic field geometry, and key components in the simulation. Both panels: The magnetic field structure can be roughly divided into four distinct zones, delineated by corresponding magenta, cyan, and green lines in both the left and right panels. The magenta line encloses the permanently closed stellar magnetosphere; the cyan line marks the boundary of opened stellar field lines that still thread the star; the green line outlines the extent of the large-scale magnetic field lines threading the disk. The region between the cyan and green lines corresponds to a reconnecting, outflow-driving zone where the interaction between stellar and disk fields powers the outflow.
III.1 Magnetic flux evolution
To quantify the opening of the magnetosphere field during simulation, Figure 3 shows the downward magnetic flux ($\hat{z}$ direction) at $z=\pm 5~\mathrm{au}$ as a function of time for the upper and lower hemispheres, respectively. Since the disk’s magnetic field is oriented upward, any downward flux at these heights must originate from the opening of the stellar magnetosphere. For reference, Figure 3 also shows the total stellar magnetospheric flux, defined as the total downward magnetic flux at the midplane. Due to magnetic reconnection of the innermost magnetospheric loops inside the star, the total stellar magnetic flux decreases slightly at the beginning of the simulation before stabilizing at approximately $8\times 10^{26}~\mathrm{G\ cm^{2}}$. The opened magnetospheric flux remains around $8\times 10^{25}~\mathrm{G\ cm^{2}}$, accounting for about 10% of the total magnetospheric flux.
@@ III.1 Magnetic flux evolution
Although the newly opened field lines within $1~\mathrm{au}$ originate from the stellar magnetosphere rather than the original disk field, they thread the disk vertically and link the dense midplane to the less dense surface layers, just as the original disk field. This connection facilitates the launching of a slower outflow at larger cylindrical radii, which we examine in the following subsection.
@@
III.2 Disk Wind from Opened Magnetospheric Field
[…]
Tu et al. (2025a) identify two components for the disk wind: an MRI-active disk wind and a laminar disk wind. In their case, the inner disk wind is MRI-unstable, leading to a turbulent outflow, whereas the outer disk wind is MRI-stable, resulting in a laminar wind. The MRI stability is measured using the dimensionless parameter (Tu et al., 2025a)
@@ III.2 Disk Wind from Opened Magnetospheric Field
where $H=R/10$ is the disk scale height and $\lambda$ is the most unstable MRI wavelength given by (Zhu and Stone, 2018)
$\lambda=2\pi\sqrt{\frac{16}{15}}\frac{v_{A,p}}{\Omega}.$ (11)
Here, $v_{A,p}=\sqrt{(B_{r}^{2}+B_{\theta}^{2})/4\pi\rho}$ is the poloidal Alfvén speed [^2] Strictly speaking, the Alfvén speed should be computed with only the laminar component of the magnetic field. Here, we use the azimuthally-averaged quantities to yield an approximation to the laminar component. and $\Omega$ is the local orbital frequency. Following Tu et al. (2025a), we choose the range $0.1<q<10$ to denote where MRI is most efficient.
Figure 4(c) shows this $q$ parameter, illustrating a zone between the jet and the disk wind in both hemispheres where $q$ is within the optimal range. The disk wind can thus be divided into the MRI-active disk wind at a small cylindrical radius and the MRI-stable disk wind at a larger cylindrical radius, particularly along the upper hemisphere. However, similar to Tu et al. (2025a), the MRI-active disk wind eventually merges with the laminar disk wind at high altitudes, where MRI can no longer be sustained. As a result, at these altitudes, the primary distinction is between the jet and the disk wind.
@@
measuring the ratios between the thermal and ram pressures and the magnetic pressure. The plasma-$\beta<1$ in the disk wind indicates that thermal energy is negligible compared to magnetic energy. Plasma-$\beta$ is also comparable to or less than 1 in the inner disk region ($\lesssim$ 1 au), where the rather strong magnetic field drives highly variable, fast (often transonic) accretion streams in the disk and its elevated envelope, as indicated by the blue regions in the $v_{z}^{p}$ map (Figure 4b), particularly within a radius of $\sim 1$ au; the accretion streams can be seen even more clearly in a zoom-in map of radial velocity $v_{r}$ (not shown, available upon request). Kinetic-$\beta>1$ in the disk indicates that kinetic energy (primarily from rotation) dominates the magnetic energy.
@@
While the magnetic force generally points in the $\hat{R}$ direction (towards the axis), its magnitude is insufficient to counteract the outward centrifugal force. This imbalance leads to an increasing opening angle of the disk wind.
@@
and following Tu et al. (2025a), we define the projected effective magnetic pressure and tension forces in cylindrical coordinates by removing the $B_{z}\frac{dB_{z}}{dz}$ term that does not contribute to overall acceleration:
@@
It is evident that magnetic pressure drives the disk wind outflow, not the magnetic tension force. This result is consistent with the geometric argument and conclusion in Tu et al. (2025a) that only magnetic pressure can drive an outflow when open magnetic field lines thread a weakly magnetized disk.
Figure 4: Overview of the outflow, particularly the disk wind at a large cylindrical radius surrounding the jet. Panel (a) and (b) show the density and projected $\hat{z}-$direction velocity ($v_{z}^{p}$, equ. 9) respectively, overplotted by the azimuthal averaged fast-magnetosonic surface (the thick solid line); panel (c) shows the normalized MRI-turbulence wavelength (equ. 10; panel (d) and (e) show the plasma-$\beta$ and (equ. 12) kinetic-$\beta$ (equ. 13) respectively. The lower panels compare different force components in the disk wind. Panel (f) shows the ratio between the cylindrical-$\hat{R}$ direction magnetic force and centrifugal force. Panel (g), (h), (i), and (j) show the magnitude of the gas acceleration due to magnetic forces. The quantities are total gas acceleration in cylindrical-$\hat{R}$ direction, total magnetic acceleration in $\hat{z}-$ direction, the effective magnetic pressure component (equ. 17), and the effective magnetic tension component (equ. 18) respectively. These disk wind properties and forces show that the disk wind in this model is very similar to the disk wind in Tu et al. (2025a).
III.3 Opened Stellar Field
As discussed in Sec. III.1, part of the stellar magnetospheric field opens up. Half of the opened field threads the disk and becomes a “disk field”; the other half remains attached to the star and becomes an opened stellar field.
To illustrate the geometry of the opened stellar magnetic field, Figure 5(a) presents a density slice through the meridional plane, overlaid with magnetic field lines. On the displayed scale, the magnetic field in the polar regions above and below the star points in the $\hat{z}$ direction, opposite to the field threading the disk (e.g., Figure 4), as expected.
@@ III.3 Opened Stellar Field
Although the majority of the open stellar field does not directly contribute to jet launching, it serves as a structural backbone for the jet in the magnetically dominated region. As we demonstrate in Section III.4, the jet is primarily driven by toroidal magnetic pressure. However, a magnetic tower consisting predominantly of toroidal loops is inherently susceptible to kink instabilities (Moll et al., 2008; Huarte-Espinosa et al., 2012). The strong open stellar magnetic field in the polar region, which accounts for most of the open stellar flux, helps stabilize the jet by threading these loops along the polar axis, mitigating instabilities.
@@
Figure 5: A zoom-in view focusing on the permanently closed stellar magnetosphere (around the midplane) and the opened stellar magnetic field threading the star (in the polar region above and below the star). Panel (a) shows the real density traced by the “passive scalars”, highlighting the magnetospheric accretion streams. The sphere in the middle is the inner boundary $r_{\mathrm{fix}}$, where a star resides (see sec. II for a more detailed description of the setup). Panels (b) and (c) show the $B_{z}$ field strength and $v_{z}^{p}$ (equ. 9) respectively, illustrating the magnetic and dynamical properties of the region close to the star. Panels (d) and (e) show the plasma-$\beta$ (equ. 12) and kinetic-$\beta$ (equ. 13) respectively, indicating the magnetosphere and the polar cavity are both magnetically dominated.
III.4 Magnetospheric jet
@@
The jet is launched from the interaction between the star, its magnetosphere, and the surrounding disk, and propagates in the low-density polar cavity. Figure 6 outlines the jet’s location and key characteristics. At lower altitudes ($\lesssim 0.5~\mathrm{au}$), the jet is confined between an infall region along the opened stellar field near the polar axis (approximately delineated by the magenta line) and the disk wind/atmosphere at larger cylindrical radii (indicated by the green dashed line). At higher altitudes, the jet expands to fill the polar cavity, as shown in Figure 4[b]. The mass loss rate of the jet remains around $2\times 10^{-9}M_{\odot}/\mathrm{yr}$ in our model (see Fig. 12 below).
To assess the dynamic importance of the magnetic field, we show in Figure 6(c) and (d) the poloidal fast magnetosonic Mach number ($M_{\mathrm{fast}}^{\mathrm{pol}}=v_{\mathrm{pol}}/v_{\mathrm{fast}}$) and the poloidal Alfvén Mach number ($M_{\mathrm{Ap}}^{\mathrm{pol}}=v_{\mathrm{pol}}/v_{\mathrm{A,pol}}$), respectively. The poloidal Alfvén speed $v_{\mathrm{A,pol}}$ is computed using only the poloidal component of the magnetic field. Both $M_{\mathrm{fast}}^{\mathrm{pol}}$ and $M_{\mathrm{Ap}}^{\mathrm{pol}}$ are less than 1 in the jet-launching polar region, indicating the jet launching process is magnetically-dominated, as expected. However, we note that the jet becomes kinematically dominated at large heights (Figure 4[e]) as energy is transferred from the magnetic field to the gas (see Section V.3 below for more details). The jet is well collimated at large distances (see Fig. 1 and Fig. 14e), likely primarily by lateral confinement from the denser, more slowly moving disk wind surrounding it. Internal “hoop” stresses from its toroidal magnetic field may also contribute (see Fig. 4g, which shows that the magnetic force in the cylindrical radial direction points inward toward the axis within the jet).
@@
Since both $M_{\mathrm{fast}}^{\mathrm{pol}}$ and $M_{\mathrm{Ap}}^{\mathrm{pol}}$ are below order unity in the jet-launching polar cavity, we focus on the magnetic forces to establish a foundation for our detailed examination of the jet-launching mechanism in section  IV. The lower panels of Figure 6 illustrate the magnetic forces in the jet region. Figure 6[e] shows the total magnetic force, revealing that the acceleration of the outflow occurs primarily along the surface of the vertically extended disk, which also marks the boundary of the jet-launching cavity. Above approximately 1 au in both hemispheres, acceleration becomes less coherent (with alternating regions of acceleration and deceleration [^3] These accelerations and decelerations can potentially introduce shocks in the jet, which can result in heating of the jet.), indicating that the bulk of the flow acceleration has largely completed beyond this altitude. This is confirmed by velocities in the $\hat{z}$ direction reaching several $10^{7}~\mathrm{cm~s^{-1}}$ below 1 au, with minimal subsequent increase in velocity above this height.
@@
As demonstrated in Tu et al. (2025a), in systems with ordered, large-scale magnetic fields threading a weakly magnetized disk, only magnetic pressure forces, not magnetic tension forces, can launch outflows in global models. Our simulations confirm this finding. Figure 6(f) and (g) display the effective magnetic tension acceleration (equ. 18) and effective magnetic pressure acceleration (equ. 17), respectively. Magnetic tension acceleration generally points toward the midplane, while magnetic pressure acceleration points away from it. More specifically, Figure 6(h) shows the acceleration due to toroidal magnetic pressure gradient, defined as
@@
The toroidal magnetic pressure accounts for most of the magnetic pressure-driven jet acceleration. In the following section, we will examine the mechanism generating this crucial toroidal magnetic field (Section IV).
Figure 6: Overview of the jet. Upper panels focus on the magneto-hydrodynamic properties of the jet; lower panels focus on the magnetic forces. Panels (a) and (b) show the density and projected $\hat{z}-$direction ($v_{z}^{p}$, equ. 9) respectively, showing the jet lies in the low-density polar cavity. Panels (c) and (d) show the poloidal fast Mach number and poloidal Alfvén Mach number, respectively, showing the jet is a magnetically-dominated structure. Panels (e), (f), (g), and (h) show the total magnetic force, the effective magnetic tension force (equ. 18), the effective magnetic pressure force (equ. 17), and the toroidal magnetic pressure force (equ. 19) respectively.
IV Jet launching mechanism: “Load-fire-reload”
This section presents a detailed description and analysis of the jet-launching mechanism in our model. The jet is launched by magnetic field lines connecting the star and the disk (we name them “two-legged” field lines, see Figure 2[b] for an illustration), through a cyclic process comprising three distinct phases: load, fire, and reload. Each phase corresponds to characteristic changes in the magnetic field configuration and dynamics. We emphasize here that each cycle occurs on extremely short time scales, and only their global effects can be captured. In the following subsections, we describe each step in detail and analyze how they collectively drive the jet-launching behavior observed in the simulation.
IV.1 Load: charging $B_{\phi}$ by star-disk differential rotation